# Measurement-driven quantum advantages in shallow circuits

### v2

@@ Appendix C Appendix C: Geometrically local quantum circuit cost in two-dimensional lattices
$U=\mathcal{P}\exp\left(-\mathrm{i}\int_{0}^{1}H(s)ds\right)$ (S55)
@@
$\frac{d}{ds}S(\psi(s),\ell)\leq\eta\sum_{j:O_{j}\text{ crosses }\ell}|y_{j}(s)|,$ (S59)
@@ Appendix D Appendix D: Measurement‑based multibody XY quantum reservoir
$\mathcal{R}_{\mathsf{MD}}^{(T)}(\rho)=\bigl(U_{\mathsf{MD}}(\tau)\bigr)^{K}\,\rho\,\bigl(U_{\mathsf{MD}}(\tau)\bigr)^{\dagger K}.$ (S66)
@@
$\tilde{c}_{i}^{x}:=\begin{cases}\frac{\varepsilon}{n}\,c_{i}^{x}&|A_{i}|_{\mathcal{S}}=1\,(\mathrm{mod}\,2),\\ c_{i}^{x}&|A_{i}|_{\mathcal{S}}=0\,(\mathrm{mod}\,2),\end{cases}$ (S73)
with $\varepsilon\ll 1$, $c_{i}^{x}$ being the initially sampled coefficients, $|A_{i}|_{\mathcal{S}}$ is the overlap of row $A_{i}$ with $\mathcal{S}$.
@@
$\tilde{c}_{i}^{y}:=\begin{cases}\frac{\varepsilon}{n}\,c_{i}^{y}&|A_{i}|_{\mathcal{S}}=1\,(\mathrm{mod}\,2),\\ c_{i}^{y}&|A_{i}|_{\mathcal{S}}=0\,(\mathrm{mod}\,2).\end{cases}$ (S74)
@@
so their contribution to the expectation of $O$ is negligible. Rows with an even overlap $|A_{i}|_{\mathcal{S}}=0,2$ commute with $O$ and retain their sampled strength. With these choices one Floquet period leads to
$\operatorname{Tr}\bigl[O\,\mathcal{R}_{\mathsf{MD}}^{(T)}(|\psi_{0}\rangle)\bigr]-\operatorname{Tr}\bigl[O\,\mathcal{R}_{\mathsf{MD}}^{(T)}(|\psi_{1}\rangle)\bigr]=2-\mathcal{O}\bigl(\frac{\varepsilon}{n}\bigr).$ (S76)
@@ Appendix E Appendix E: Additional numerical benchmarks for the measurement‑based quantum reservoir
The reservoir dynamics are governed by Hamiltonians with $s=16$ terms, where all coupling strengths are sampled i.i.d. from $\mathcal{N}(0,1/2)$. The system undergoes ten identical Floquet cycles. We consider two hardware-motivated layouts: heavy-
hex and square lattices [Figs. S7(a,b)]. For each layout, we sample 10 independent architecture matrices (either randomized fan-out staircases or degree-matched local edge sets), meaning each data point represents an average over 20 distinct architectures. The corresponding standard deviations are shown as error bars.

### v3

@@ Appendix C Appendix C: Geometrically local quantum circuit cost in two-dimensional lattices
$U=\mathcal{P}\exp\left(-\mathrm{i}\int_{0}^{1}\mathcal{H}(s)\,ds\right)$ (S55)
@@
$\frac{d}{ds}S(\psi(s),\ell)\leq\eta\sum_{e\in\mathcal{E}:\,O_{e}\text{ crosses }\ell}|y_{e}(s)|$ (S59)
@@ Appendix D Appendix D: Measurement‑based multibody XY quantum reservoir
$\mathcal{R}_{\mathrm{MD}}^{(T)}(\rho)=\bigl(U_{\mathsf{MD}}(\tau)\bigr)^{K}\,\rho\,\bigl(U_{\mathsf{MD}}(\tau)\bigr)^{\dagger K}.$ (S66)
@@
$\tilde{c}_{i}^{x}:=\begin{cases}\frac{\varepsilon}{n}c_{i}^{x}&|\mathbf{A}_{i}|_{\mathcal{S}}=1\,(\mathrm{mod}\,2),\\ c_{i}^{x}&|\mathbf{A}_{i}|_{\mathcal{S}}=0\,(\mathrm{mod}\,2),\end{cases}$ (S73)
with $\varepsilon\ll 1$, $c_{i}^{x}$ being the initially sampled coefficients, $|\mathbf{A}_{i}|_{\mathcal{S}}$ is the overlap of row $A_{i}$ with $\mathcal{S}$.
@@
$\tilde{c}_{i}^{y}:=\begin{cases}\frac{\varepsilon}{n}\,c_{i}^{y}&|\mathbf{A}_{i}|_{\mathcal{S}}=1\,(\mathrm{mod}\,2),\\ c_{i}^{y}&|\mathbf{A}_{i}|_{\mathcal{S}}=0\,(\mathrm{mod}\,2).\end{cases}$ (S74)
@@
so their contribution to the expectation of $O$ is negligible. Rows with an even overlap $|\mathbf{A}_{i}|_{\mathcal{S}}=0,2$ commute with $O$ and retain their sampled strength. With these choices one Floquet period leads to
[…]
@@ Appendix E Appendix E: Additional numerical benchmarks for the measurement‑based quantum reservoir
The reservoir dynamics are governed by Hamiltonians with $s=16$ terms, where all coupling strengths are sampled i.i.d. from $\mathcal{N}(0,1/2)$. The system undergoes ten identical Floquet cycles. We consider two hardware-motivated layouts: heavy-hex and square lattices [Figs. S7(a,b)]. For each layout, we sample 10 independent architecture matrices (either randomized fan-out staircases or degree-matched local edge sets), meaning each data point represents an average over 20 distinct architectures. The corresponding standard deviations are shown as error bars.